OpenStudy (anonymous):

solve by substitution or elimination: xy=8 and x+y= 6

OpenStudy (anonymous):

xy = 8 y = 8/x plug 8/x into y in the equation x+y=6. solve for x.

OpenStudy (anonymous):

would it be (-y+6)y=8

OpenStudy (anonymous):

that's a good start... no solve for y...

OpenStudy (anonymous):

*now

OpenStudy (anonymous):

actually it would be y = 8/x x+y = 6 x + (8/x) = 6 solve for x.

OpenStudy (anonymous):

then i keep getting -y^2+6y-8+0 but it isnt factoring right

OpenStudy (anonymous):

i mean =0

OpenStudy (anonymous):

what would be the solution to this problem?

OpenStudy (anonymous):

im still not getting the right answer.

OpenStudy (anonymous):

you mean: -y^2+6y-8=0 if it doesn't factor, that means you have irrational solutions... BUT this one does factor..

OpenStudy (anonymous):

you can actually write it out as \(\large y^2-6y+8=0 \) so \(\large (y-4)(y-2)=0 \)

OpenStudy (anonymous):

xy=8 and x+y= 6 y = 8/x x + (8/x) = 6 x^2 + 8 = 6x x^2 -6x + 8 = 0 (x-4)(x-2) = 0 x = 4,2 plug into the original equation to get the y values solve using the quadratic formula